### v2

# ReverseNER: A Self-Generated Example-Driven Framework for Zero-Shot Named Entity Recognition with Large Language Models

Abstract
This paper presents ReverseNER, a framework aimed at overcoming the limitations of large language models (LLMs) in zero-shot Named Entity Recognition (NER) tasks, particularly in cases where certain entity types have ambiguous boundaries. ReverseNER tackles this challenge by constructing a reliable example library with the reversed process of NER. Rather than beginning with sentences, this method uses an LLM to generate entities based on their definitions and then expands them into full sentences. During sentence generation, the LLM is guided to replicate the structure of a specific ’feature sentence’, extracted from the task sentences by clustering. This results in well-annotated sentences with clearly labeled entities, while preserving semantic and structural similarity to the task sentences. Once the example library is constructed, the method selects the most semantically similar example labels for each task sentence to support the LLM’s inference. We also propose an entity-level self-consistency scoring mechanism to improve NER performance with LLMs. Experiments show that ReverseNER significantly outperforms traditional zero-shot NER with LLMs and surpasses several few-shot methods, marking a notable improvement in NER for domains with limited labeled data.
keywords:
[…]
1 Introduction
The advent of Large Language Models (LLMs) has marked a significant technological leap for Natural Language Processing (NLP). Owing to extensive pre-training and parameter optimization, LLMs such as GPT-3 [2] and BERT (Bidirectional Encoder Representations from Transformers) [6] have demonstrated exceptional performance across various NLP tasks. By capturing semantics, syntax, and contextual relationships from vast corpora, these models have not only achieved breakthroughs in traditional tasks like machine translation [14], relation extraction [17], and Named Entity Recognition (NER) [16], but also exhibited strong generalization capabilities in generative tasks such as text generation [12] and dialogue generation.
Despite the impressive performance of LLMs in multiple tasks, they still face challenges in zero-shot NER, which requires models to correctly identify entity types without having seen any previously annotated data. In contrast, traditional NER methods typically rely on a large amount of annotated data for supervised learning [9]. Although LLMs possess some transfer learning capabilities, when enabling them to complete certain NER tasks without annotations, they still constantly struggle with complex tasks due to insufficient generalization abilities and imprecise understanding of domain-specific entities. These shortcomings arise from the model’s dependence on the diversity and richness of the training corpora for recognizing specific entity contexts, and when faced with new entity types or contexts, the models often struggle to match the entity categories accurately.
To address the limitations of LLMs in zero-shot NER tasks, we propose a novel method named ReverseNER, which constructs an example library by reversing the NER process. This method first utilizes a pre-trained BERT model to calculate the similarity between task sentences and clusters the sentences based on this similarity. The LLM uses the central sentences of each cluster to generate similar sentences and construct a high-quality example library in conjunction with entity types. Next, the method calculates the cosine similarity between the task sentences and the sentences in the example library, selects the closest examples from the library, and adds them to the prompt when doing actual NER tasks. Compared to traditional zero-shot approaches using LLMs, our method provides LLMs with accurate, pre-generated examples that are semantically similar to the sentences requiring recognition, significantly enhancing NER task performance.
The contributions of this paper are as follows: (1) This paper proposes a method called ReverseNER for large models to perform zero-shot NER. (2) ReverseNER generates accurate and task-relevant examples by combining task sentence syntax and entity types to reverse the traditional NER process. (3) Evaluation results demonstrate that our method in this paper has achieved significant improvements on three datasets.
2 Related work
2.1 Named Entity Recognition
Named Entity Recognition is a fundamental task in Natural Language Processing, aimed at detecting named entities from text and categorizing them into predefined entity types. Traditional research approaches frame the NER task as a sequence labeling problem. For instance, Hammerton et al. [7] utilizes unidirectional LSTMs to obtain token-level representations, which are then fed into a softmax classifier for entity classification; Collobert et al. [3] employs CNNs to embed each input word and used Conditional Random Fields (CRFs) to decode each embedding into entity labels; Devlin et al. [6] acquires token-level representations through BERT for classification purposes; Sarzynska-Wawer et al. [13] enhances the quality of word representations by using large-scale pre-trained models; Li et al. [10] formulates the NER task as a Machine Reading Comprehension (MRC) task and further optimizes the performance of the MRC model using the dice loss function.
2.2 Applications of LLM in NER
In recent years, numerous studies have explored the application of LLMs to NER tasks. One of the earliest advancements is introduced by Devlin et al. [6] with the BERT model, where fine-tuning on specific NER datasets significantly improves performance due to BERT’s ability to capture bidirectional contextual information, which is crucial for understanding entity boundaries in text.
Following this, RoBERTa, an optimized version of BERT, is introduced by Liu et al.[11], which demonstrates further improvements in NER by training on larger datasets with better hyperparameter tuning, dynamic masking, and longer training duration. RoBERTa outperformed BERT on several NER benchmarks.
Additionally, there has been an exploration of transforming NER into a generative task. Yan et al. proposed using GPT models [2] for NER by generating sentences with labeled entities instead of relying on the traditional sequence labeling format. This shift toward a generation-based approach allows for more flexible zero-shot learning settings, particularly useful in scenarios with minimal labeled data [19].
Prompt-based learning methods also gained traction for NER, especially in few-shot and zero-shot scenarios. Cui et al[4]. investigated how cloze questions and prompts could be used to leverage LLMs’ pretraining capabilities for recognizing entities with minimal task-specific training. This method shows promise in extracting entities with little to no labeled examples.
Most recently, Xie et al[18]. introduced a novel self-improving framework for zero-shot NER. Their approach utilizes an LLM to generate self-annotated datasets from an unlabeled corpus. The system then selects reliable pseudo-labeled data through self-consistency (SC) filtering strategies, and finally, this refined dataset is used for in-context learning (ICL) during inference. This self-improving strategy significantly enhances zero-shot performance on NER tasks across several benchmarks, including CoNLL03 [15] and WikiGold [1].
3 Method
Our insight lies in the idea that providing reference examples can significantly enhance the performance of LLMs in NER tasks, making it crucial to ensure that these examples are both accurate and precisely aligned with the specific requirements and complexities of the task. To achieve this, we propose the ReverseNER method, which constructs an example library by reversing the conventional NER process: first generating the entities, and subsequently prompting the language model to generate sentences that incorporate these entities. During this sentence generation phase, a subset of sentences from the task set will be provided as structural references to guide the model. This example library aims to further enhance the model’s performance in subsequent NER tasks by offering task-specific, high-quality examples.
The primary workflow, shown in Figure 1, has three main stages: extracting feature sentences from task sets, constructing an example library with LLM, and for each task sentence, selecting the most semantically similar example labels and incorporating them into the prompt as examples.
Figure 1: Overview of the ReverseNER Process.
3.1 Feature Sentence Extraction
3.1.1 Sentence Embedding Using BERT
The BERT model is employed to map sentences from the task set into high-dimensional vector representations. BERT is a pre-trained model based on the Transformer architecture, which effectively captures contextual information through a bidirectional encoder, generating deep, high-dimensional semantic embeddings for each sentence [6]. Compared to traditional methods such as the Bag-of-Words model and TF-IDF, BERT captures the semantic similarity between texts more effectively. Specifically, given a sentence $S=[w_{1},w_{2},\ldots,w_{n}]$ composed of words $w_{i}$, BERT generates the embedding vector of the sentence denoted as $v_{s}=f(H)=f(\text{BERT}(S))$, where $f$ denotes a further processing function of the BERT output, and $H$ denotes the hidden state matrix generated by BERT.
These generated embeddings enable the clustering of sentences, which in turn facilitates the identification of representative sentences that can be used to generate high-quality reference examples.
3.1.2 Task Set Clustering
After obtaining sentence embeddings with BERT, the K-Medoids clustering algorithm is applied for sentence clustering. Unlike K-Means, K-Medoids select actual data points as cluster centers. We select the sentences corresponding to these cluster centers as feature sentences. The specific steps of the K-Medoids clustering algorithm are as follows[8]:
  • Initialize Medoids: Randomly select $k$ embeddings from the set of embedding vectors as the initial medoids.
  • Assign Embeddings to Medoids: Measure the distance between sentence embeddings using cosine similarity. For each embedding $v_{i}$, identify its nearest medoid $c_{j}$ and assign $v_{i}$ to the cluster $C_{j}$ corresponding to $c_{j}$.
  • Update Medoids: For each cluster $C_{j}$, find the embedding that minimizes the total distance to all other embeddings in the cluster and designate it as the new medoid, formulated as:
    $c_{j}=\arg\min_{v\in C_{j}}\sum_{v_{i}\in C_{j}}d(v,v_{i})$
    In other words, extract a representative embedding within the cluster such that the total distance between this embedding and all other embeddings in the cluster is minimized.
  • Iterate: Repeat the steps of assigning embeddings and updating medoids until the cluster centers no longer change or reach the maximum iterations.
The use of K-Medoids facilitates the extraction of feature sentences that are both representative and diverse. By selecting the cluster centers that capture different aspects of the dataset, we can ensure that the example library includes a wide range of sentence structures and contexts. Such diversity is crucial for the LLMs to handle a variety of linguistic patterns and entities, ultimately leading to improved generalization and performance.
3.2 Example Library Construction
NER algorithms take task sentences as the input and output recognized entities. However, especially in zero-shot scenarios, these recognized results do not qualify as gold labels due to the imperfect semantic understanding of the models, and consequently, example libraries constructed from these results may lack reliability. On the contrary, if the generation algorithms take entities as input and output a sentence that includes these entities, as the reverse process of NER, the result will be much more reliable. Inspired by this, we present a novel method to construct the example library based on such a reverse process.
Specifically, the example library generation in the ReverseNER method involves two steps using LLMs: first, generating a set of entities based on their definitions, then expanding these entities into sentences.
3.2.1 Entity Vocabulary Generation
The entity types to be identified are first provided to an LLM, which then generates a series of specific entities based on each type’s definition. The LLM is instructed to ensure diversity in the generated entities. The dialogue demonstration of the entity vocabulary generation step is shown as follows (for simplicity, entity definitions are omitted):
Prompt Here is an entity type set: [PER, LOC, ORG]. (Omitted: Definitions of each entity type) Please imagine a list of at least 3 diverse words for each entity type in the set.
LLM PER: John Smith, Marie Curie, Albert Einstein LOC: Paris, Amazon Rainforest, Mount Everest ORG: United Nations, Red Cross, NASA
3.2.2 Example Label Generation
After obtaining these entity nouns, for each feature sentence extracted in Section 3.1, the LLM is employed to generate a set of new sentences according to the original sentence’s structure and format. These newly generated sentences should maintain structural consistency with the original sentences and ensure the inclusion of the previously generated entity nouns. Through this method, example labels i.e., sentences with their entities labeled, are produced, ultimately constructing a fully annotated example library to support subsequent NER tasks. The dialogue demonstration of the entity vocabulary generation step is shown as follows:
Prompt Here is an entity type set: [PER, LOC, ORG]. Construct a sentence that contains entities from the following words: PER: John Smith, Marie Curie, Albert Einstein LOC: Paris, Amazon Rainforest, Mount Everest ORG: United Nations, Red Cross, NASA Output the sentence along with the utilized entities. Here is a reference sentence, try to mimic its style and content: John and Sarah work for UNICEF in New York.
LLM Sentence: John Smith and Marie Curie work for NASA in Paris. Entities: [{"John Smith":"PER"}, {"Marie Curie":"PER"}, {"NASA":"ORG"}, {"Paris":"LOC"}]
As shown in the dialogue above, the example labels generated through the reversed NER process may not always be semantically realistic, but they ensure reliability as valid example labels for NER tasks.
3.3 Example-Driven Inference
3.3.1 Example Labels Selection
To identify the most relevant example labels for each task sentence, a similarity matrix must be calculated between the sentences in the example library and those in the task set. To achieve this, we first use the BERT model to map each sentence in the example library into high-dimensional embedding vectors. Then, for each sentence in the task set, the cosine similarity between it and every sentence in the example library is calculated, resulting in a similarity matrix of size $N\times M$, where $N$ denotes the number of sentences in the task set, and $M$ denotes the number of sentences in the example library.
Based on the similarity matrix, we select the most relevant example labels to include in the prompt provided to the LLM during entity recognition.
3.3.2 Inference with LLMs
The constructed prompt to perform the NER task on a certain task sentence includes the following four elements: the entity type set to be recognized, definitions of each entity type, selected example labels, and the task sentence itself. The dialogue demonstration of inference is shown as follows:
Prompt Perform NER task on this entity type set: [PER, LOC, ORG]. (Omitted: Definitions of each entity type) Example 1: Input Sentence: John Smith and Marie Curie work for NASA in Paris. Output: [{"John Smith":"PER"}, {"Marie Curie":"PER"}, {"NASA":"ORG"}, {"Paris":"LOC"}] (Omitted: Example 2 to Example 5) Task Input Sentence: Lily and Mark collaborate at the World Health Organization in Geneva. Your output:
LLM [{"Lily":"PER"}, {"Mark":"PER"}, {"World Health Organization":"ORG"}, {"Geneva":"LOC"}]
3.4 Entity-Level Self-Consistency Scoring
To mitigate the impact of LLM hallucinations, we perform multiple reasoning attempt rounds on the same sentence to be recognized, and introduce the self-consistency (SC) scoring mechanism. Different from the traditional Response-Level SC methods [wang2022self], we propose the Entity-Level SC Scoring method.
The occurrences of each inferred entity type across all reasoning attempts are counted, which is used as the self-consistency (SC) score for each entity. For each reasoning attempt, the score is calculated as the average SC score of all entities inferred in that response. The average SC score for each reasoning is formulated as:
$SC_{average}=\frac{1}{n}\sum_{i=1}^{n}SC(e_{i})$
where $n$ denotes the number of entities in the reasoning attempt, and $SC(e_{i})$ denotes the self-consistency score of entity $e_{i}$, which is the total number of times entity $e_{i}$ appears across all rounds of attempts.
Additionally, to handle cases where the correct output does not contain any entities, a majority-based rule is introduced: if more than half of the predictions indicate no entity, the final output is set to "no entity."
For each sentence in the task set, the self-consistency scoring method is applied during testing, and the result with the highest self-consistency score is selected as the final output.
4 Experiments
In this section, we experiment on four public benchmarks to evaluate the ReverseNER method in comparison to multiple baselines.
4.1 Setup
LLM
We utilize the GPT-4o mini model as the LLM to generate examples and perform the NER task to evaluate our algorithm. During the self-consistency scoring process, the temperature parameter is set to 0.8, and the prediction process is repeated five times.
Example Library Generation
For English texts, we utilize the BERT-base-uncased model [6] to derive sentence embeddings, while for Chinese texts, we use the BERT-base-Chinese model [6]. Sentences from the input set are clustered into 10 categories, and for each cluster’s central sentence i.e., the feature sentences, the LLM generates 3 examples by mimicking the sentence structure. Thus, these 30 example labels constitute the entire example library for a certain dataset.
Datasets
We use the following four datasets to evaluate the performance of the NER task. For English texts, we utilize the CoNLL03 dataset [15] and WikiGold dataset [1]. For Chinese texts, we use the People’s Daily dataset [5] and have constructed a novel dataset called GovAff, which consists of sentences pertaining to routine administrative affairs. The GovAff dataset encompasses four types of entities: person, government agency, corporation, and miscellaneous.
Baselines and Comparison
The performance achieved with the vanilla model, where no examples are generated or provided, and only the task sentence and entity types are provided to the model, is recorded as the baseline. Additionally, we conduct two few-shot NER experiments using 10 examples with gold labels and compare their results with the proposed method in this study.
4.2 Overall Performance
The results presented in Table 1 underscore the F1 score of the zero-shot ReverseNER method when compared to other methods. Notably, the zero-shot ReverseNER with self-consistency scoring attains the highest average F1 score of 79.10 across the four evaluated datasets, outperforming both the Vanilla zero-shot and few-shot baselines.
Relative to the Vanilla zero-shot baseline, the zero-shot ReverseNER with SC demonstrates a substantial enhancement in the average F1 score (79.10 vs. 71.22), reflecting a significant improvement in the model’s capacity to generalize to unseen data without the need for additional gold labels provided in advance. Furthermore, in comparison to the few-shot methods, both with and without SC, the zero-shot ReverseNER with SC exhibits superior performance, surpassing the few-shot SC approach (77.25 average) by 1.85 points.
Table 1: F1 Score of ReverseNER method and its comparisons.
Method CoNLL03 WikiGold People’s Daily GovAff Average
Vanilla 71.69 72.39 69.56 71.23 71.22
Few-shot, w/o SC 72.71 75.86 71.23 83.40 75.80
Few-shot, SC 74.21 76.76 73.89 84.14 77.25
Zero-shot ReverseNER, w/o SC 76.19 76.98 75.46 82.46 77.77
Zero-shot ReverseNER, SC 77.78 78.47 76.62 83.53 79.10
4.3 Analysis
4.3.1 Detailed Performance
Taking the WikiGold dataset as a case study, we conduct a detailed analysis of the recognition accuracy for each label. Results are shown in Table 2. The WikiGold dataset comprises four types of entities: PER, ORG, LOC, and MISC, denoting persons, organizations, locations, and miscellaneous entities, respectively. Notably, the MISC label is more ambiguous in definition, covering entities that do not fall into the aforementioned categories but still qualify as proper nouns. Consequently, the recognition accuracy for this label tends to be suboptimal.
Table 2: Detailed NER Recognition Results. Methods with the best Precision for each metric are indicated with underlining, those with the best Recall are italicized, and those with the highest F1 score are highlighted in bold.
Method Metric Entities Overall
PER ORG LOC MISC
Vanilla Precision 87.14 70.91 83.33 32.73 63.97
Recall 85.98 72.84 85.63 71.62 79.69
F1 86.56 71.86 84.46 44.93 70.97
Zero-shot ReverseNER, w/o SC Precision 90.93 77.79 87.75 47.86 74.96
Recall 84.41 72.05 85.42 72.50 79.11
F1 87.39 74.81 86.57 57.66 76.98
Zero-shot ReverseNER, SC Precision 92.16 76.88 86.89 54.39 77.66
Recall 85.65 73.30 85.12 70.15 79.29
F1 88.79 75.04 85.99 61.27 78.47
By examining the Precision, Recall, and F1 scores of each label across different methods, we observe that the ReverseNER with SC voting consistently achieves the highest performance in terms of Precision and F1 scores for most labels, followed by the ReverseNER without SC voting. The Vanilla method, however, frequently achieves higher Recall, particularly for PER and LOC types. This phenomenon can be attributed to two primary factors: (1) The conservative strategy employed by the LLM when generating examples, where uncertain terms that may not belong to the current entity label are often excluded, leads to higher Precision but potentially lower Recall; (2) The SC voting algorithm tends to penalize outputs containing entities that are not consistently present across all outputs, even if they are correct in some cases, which further boosts Precision while reducing Recall. Consequently, these strategies improve the F1 score, particularly for labels with less clearly defined boundaries such as MISC categories, as they achieve a better balance between Precision and Recall.
4.3.2 Impact of hyperparameters
Our ReverseNER method involves two hyperparameters: the number of clusters used during feature sentence extraction and the number of example labels that the LLM needs to construct per cluster. The product of these two hyperparameters determines the total number of labels in the example library. To evaluate the algorithm’s stability under different parameter settings, we do not enable the self-consistency scoring mechanism and run each hyperparameter combination 20 times, calculating the mean and standard deviation.
Taking the WikiGold dataset as an example, Table 3 illustrates the performance for each set of hyperparameter combinations. From the table, we can observe that keeping the total number of labels in the example library constant while increasing the number of clusters and decreasing the number of labels constructed per cluster leads to decreased standard deviation, standing for increased stability. However, for the F1 score, there is a peak point; either increasing or decreasing the number of clusters beyond this point negatively impacts the results.
The reason for this phenomenon may be that when the number of clusters is small, more labels need to be constructed for each cluster, resulting in a large number of labels with similar sentence structures in the example library. This makes it difficult for the LLM to receive diverse and contextually appropriate examples during each inferring, thereby reducing both the overall F1 score and stability. Conversely, when the number of clusters is too large, the sentence structures in the example library become more varied, but it also becomes more challenging to select the most contextually relevant examples for the current task sentence, leading to a slight decrease in the F1 score.
Notably, as it is shown in Figure 2, when the number of clusters is fixed, simply increasing the number of labels constructed for each cluster is unlikely to bring performance improvement and may instead lead to a decrease in F1 score stability.
Figure 2: F1 Score on WikiGold when the number of clusters is fixed at 10.
Table 3: F1 Score on WikiGold with different sets of hyperparameters, without self-consistency scoring
Clusters Labels per Cluster F1 Score
1 30 75.60 $\pm$ 2.15
2 15 75.72 $\pm$ 1.78
3 10 76.02 $\pm$ 1.18
5 6 76.07 $\pm$ 1.08
6 5 76.71 $\pm$ 0.78
10 3 76.98 $\pm$ 0.71
15 2 76.28 $\pm$ 0.59
30 1 76.19 $\pm$ 0.57
1 60 74.23 $\pm$ 3.60
2 30 75.04 $\pm$ 1.43
3 20 75.31 $\pm$ 1.29
4 15 75.43 $\pm$ 1.21
5 12 75.85 $\pm$ 1.18
6 10 75.52 $\pm$ 0.96
10 6 76.03 $\pm$ 0.89
12 5 76.14 $\pm$ 0.95
15 4 76.63 $\pm$ 0.86
20 3 76.93 $\pm$ 0.78
30 2 75.90 $\pm$ 0.73
60 1 76.28 $\pm$ 0.54
4.3.3 Effect of Entity-Level Self-Consistency Scoring
The comparison presented in Table 4 elucidates the performance differences between the Entity-Level Self-Consistency (SC) and the traditional Response-Level SC methods, which operate by selecting the most frequently occurring result from multiple iterations, across various datasets, evaluated using the F1 score metric. On the CoNLL03, WikiGold, and GovAff datasets, the Entity-Level SC method performs slightly better than the Response-Level SC method. In contrast, the Response-Level SC method outperforms the Entity-Level SC on the People’s Daily dataset.
This isolated underperformance for the Entity-Level SC can be attributed to the fact that there are only three clearly defined types of entities in the People’s Daily Dataset: PER, LOC, and ORG, without the MISC type in other datasets, which does not have a well-defined boundary. The Entity-Level SC strategy emphasizes trading Recall for Precision to enhance the F1 score, but this approach is less effective in datasets where entity definitions are more straightforward, resulting in lower performance compared to the traditional Response-Level SC method.
Table 4: F1 Score comparison on different SC methods.
Dataset Entity-Level SC Response-Level SC
CoNLL03 77.78 77.41
WikiGold 78.47 78.00
People’s Daily 76.62 77.72
GovAff 83.53 83.34
5 Conclusion
In conclusion, ReverseNER provides a sophisticated methodology for addressing zero-shot NER by leveraging self-generated examples within an example-driven framework. By constructing a labeled example library and employing self-consistency scoring, ReverseNER effectively mitigates the challenges associated with NER in the absence of annotated datasets. Experimental evaluations indicate substantial improvements over conventional zero-shot and few-shot NER approaches, underscoring its potential for enhancing entity recognition capabilities. Future research endeavors will focus on optimizing computational efficiency and broadening the applicability of ReverseNER to encompass more diverse NLP tasks.
6 Limitations
We acknowledge the following limitations of this study:
  • The proposed method may lead to a slight decline in Recall due to trade-offs between Precision and Recall.
  • The self-consistency scoring mechanism increases the volume of reasoning attempts, which may escalate runtime and incur additional costs, raising scalability concerns in resource-limited settings.
7 Acknowledgment
This work is funded by the Xi’an Jiaotong University-China Mobile Communications Group Co., Ltd. Digital Government Joint Institute.
Appendix A Actual Prompts for Constructing the Example Library
Taking the WikiGold dataset as an example, the prompt used for generating entity vocabularies of this dataset is illustrated in Listing A1. We repeat this prompt when constructing example labels for each cluster.
The prompt for constructing example labels in the example library is shown in Listing A2. The entity vocabularies it uses are derived from the results of the corresponding entity vocabulary generation prompt. The reference sentence provided at the end corresponds to the sentence located at the center of the respective cluster. We repeat this prompt when constructing example labels for each cluster.
Here is an entity type set: [PER, ORG, LOC, MISC].
Here are the explanations of each entity from the label set:
PER means Person. Definition: Denotes individual people or fictional characters. This includes full names, nicknames, and titles when they are part of the name.
ORG means Organization. Definition: Represents groups of people that are identified by a particular name. This includes companies, institutions, government bodies, agencies, and other formal organizations.
LOC means Location. Definition: Refers to geographical entities such as countries, cities, landmarks, mountains, rivers, and any other physical locations.
MISC means Miscellaneous. Definition: Covers entities that do not fall into the above categories but are still proper nouns. This includes nationalities, religions, events, languages, works of art, and other entities.
Please imagine a list of at least 6 diverse words for each entity type in the set. The output must be a JSON object, where keys are the entity types and values are lists of words. Here is the output JSON structure example you must follow:
"PER": ["PER_1", "PER_2", "PER_3"],
"ORG": ["ORG_1", "ORG_2", "ORG_3"],
"LOC": ["LOC_1", "LOC_2", "LOC_3"],
"MISC": ["MISC_1", "MISC_2", "MISC_3"]
Listing 1: Vocabulary Generation Prompt
Here is an entity label set: [PER, ORG, LOC, MISC], do as follows:
Make 3 English sentences with diverse styles and word orders.
Each sentence must contain one, two, or three entities from the following words:
PER: Leonardo da Vinci, Nelson Mandela, Marie Curie, Frida Kahlo, Sherlock Holmes, Malala Yousafzai
ORG: NASA, Amnesty International, Harvard University, World Wildlife Fund, Microsoft, International Monetary Fund
LOC: Tokyo, Grand Canyon, Great Wall of China, Nile River, Mount Kilimanjaro, Sydney
MISC: Jazz, Buddhism, Renaissance, Spanish, The Great Gatsby, Nobel Prize
If your generated sentences include entities that may belong to a certain type of entity label set but are not shown in the word list, mark them as well in the output JSON.
Output format:
"text": "Your sentence here.",
"entities": [
{"entity_text": "the PER entity","entity_label": "PER"},
{"entity_text": "the ORG entity","entity_label": "ORG"},
{"entity_text": "the LOC entity","entity_label": "LOC"},
{"entity_text": "the MISC entity","entity_label": "MISC"}
Here is a sentence for reference. You can reference its style and content, but ensure the generated sentence covers different topics and scenarios
Reference Sentence: He died in Hollywood, California.
Listing 2: Label Generation Prompt
Appendix B Actual Prompts for Performing NER on Each Task Input
The prompt used for inferring a specific input sentence from the WikiGold dataset is shown in Listing B3. The provided examples are selected from the constructed example library based on cosine similarity.
Perform English NER task in the following entity type set: [PER, ORG, LOC, MISC]
Here are the explanations of each entity from the label set:
PER means Person. Definition: Denotes individual people or fictional characters. This includes full names, nicknames, and titles when they are part of the name.
ORG means Organization. Definition: Represents groups of people that are identified by a particular name. This includes companies, institutions, government bodies, agencies, and other formal organizations.
LOC means Location. Definition: Refers to geographical entities such as countries, cities, landmarks, mountains, rivers, and any other physical locations.
MISC means Miscellaneous. Definition: Covers entities that do not fall into the above categories but are still proper nouns. This includes nationalities, religions, events, languages, works of art, and other entities.
Output format: [{"entity 1 text": "entity 1 type"},{"entity 2 text": "entity 2 type"}]
Example 1: Input sentence: Marie Curie was awarded the Nobel Prize for her groundbreaking research in radioactivity. Output: [{"Marie Curie": "PER"}, {"Nobel Prize": "MISC"}]
Example 2: Input sentence: The World Cup is celebrated globally, bringing nations like Brazil and Germany together. Output: [{"World Cup": "MISC"}, {"Brazil": "LOC"}, {"Germany": "LOC"}]
Example 3: Input sentence: In 2019, Greenpeace launched a campaign focused on the Great Wall of China to raise awareness about environmental issues. Output: [{"Greenpeace": "ORG"}, {"Great Wall of China": "LOC"}]
Example 4: Input sentence: Nelson Mandelas advocacy for peace was often celebrated during Diwali festivities around the world. Output: [{"Nelson Mandela": "PER"}, {"Diwali": "MISC"}]
Example 5: Input sentence: Marie Curies groundbreaking research in radiation opened new doors for the field of science at Harvard University. Output: [{"Marie Curie": "PER"}, {"Harvard University": "ORG"}]
Test Input Sentence: Another Nobel laureate scientist there was Richard J. Roberts. Output:
Listing 3: Final Recognition Prompt